a rectangle is inscribed under the curve y=2^-x with its base along the positive x-axis. Find the dimensions of the rectangle with the largest area


1
Expert's answer
2022-04-21T16:23:26-0400

Define:

"x" - length of the rectangle "(x>0)"

"2^{-x}" - height of the rectangle "(x>0)"

Area of the rectangle:

"S(x)=x\cdot2^{-x}"

"S’(x)=2^{-x}-x\cdot2^{-x}\cdot\ln2=0"

"x=\frac {1}{\ln2}"

"\displaystyle S’(\frac{1}{2\ln2})=2^{-\frac{1}{2\ln2}}(1-\frac{1}{2\ln2}\cdot\ln2)>0"

"\displaystyle S’(\frac{2}{\ln2})=2^{-\frac{2}{\ln2}}(1-\frac{2}{\ln2}\cdot\ln2)<0"

So "x=\frac {1}{\ln2}" corresponds to maximum of the function "S(x)" .

Answer:

"\frac {1}{\ln2}" - length of the rectangle

"2^{-\frac {1}{\ln2}}" - height of the rectangle